OpenStudy (anonymous):

Simplify the following expression: 1/ (\sqrt[3]{x^8}) That is a cube root with the argument x^8 in the denominator. ATTEMPT: Is it not: 1/ (x^2*\sqrt[3]{x^2}) ?

OpenStudy (anonymous):

\[\frac{1}{\sqrt[3]{x^8}}=\frac{1}{x^{8/3}}=x^{-8/3}\]